### v1

# HiFiSeg: High-Frequency Information Enhanced Polyp Segmentation with Global-Local Vision Transformer

Abstract
Numerous studies have demonstrated the strong performance of Vision Transformer (ViT)-based methods across various computer vision tasks. However, ViT models often struggle to effectively capture high-frequency components in images, which are crucial for detecting small targets and preserving edge details, especially in complex scenarios. This limitation is particularly challenging in colon polyp segmentation, where polyps exhibit significant variability in structure, texture, and shape. High-frequency information, such as boundary details, is essential for achieving precise semantic segmentation in this context. To address these challenges, we propose HiFiSeg, a novel network for colon polyp segmentation that enhances high-frequency information processing through a global-local vision transformer framework. HiFiSeg leverages the pyramid vision transformer (PVT) as its encoder and introduces two key modules: the global-local interaction module (GLIM) and the selective aggregation module (SAM). GLIM employs a parallel structure to fuse global and local information at multiple scales, effectively capturing fine-grained features. SAM selectively integrates boundary details from low-level features with semantic information from high-level features, significantly improving the model’s ability to accurately detect and segment polyps. Extensive experiments on five widely recognized benchmark datasets demonstrate the effectiveness of HiFiSeg for polyp segmentation. Notably, the mDice scores on the challenging CVC-ColonDB and ETIS datasets reached 0.826 and 0.822, respectively, underscoring the superior performance of HiFiSeg in handling the specific complexities of this task.
Index Terms:
polyp segmentation, colonoscopy, deep learning, pyramid vision transformer
I Introduction
Polyps are abnormal growths in the colon and rectum that protrude above the surrounding intestinal mucosa. Colorectal cancer often develops from colonic polyps, particularly adenomatous ones. Therefore, early detection and removal of these polyps are crucial for preventing the progression to colorectal cancer. Colonoscopy is the gold standard in clinical practice for detecting colorectal lesions [1]. However, manual annotation of polyps is not only labor-intensive but also prone to errors. This highlights an increasing demand for automated and precise image segmentation methods to assist doctors in diagnosis.
Deep learning algorithms represented by CNNs have achieved remarkable success in medical image applications such as cardiac, skin lesion, and polyp segmentation. In particular, fully convolutional networks (FCNs), such as UNet [2], SegNet [3], and DeepLab [4], have become the predominant methods in this field. However, due to the limited receptive fields of CNNs, these CNN-based methods have limitations in capturing long-range relationships and global context, which restricts their representational capabilities in capturing shape and structural information crucial for medical image segmentation.
Transformer [5] reflects complex spatial transformations and long-range feature dependencies through the multi-head self-attention (MHSA) mechanism. It has achieved significant success in the field of natural language processing (NLP). In recent years, the Vision Transformer (ViT) [6] has been proposed to apply the Transformer structure to vision tasks, aiming to overcome the limitations of CNNs in image recognition tasks. Unfortunately, despite the ability of Transformers to capture global dependencies, they struggle with image locality and translational invariance, which affects the accurate segmentation of small targets and boundaries. Therefore, constructing a model that retains both local and global features while preserving their respective characteristics remains a challenging task. Recent works have introduced hybrid architectures of Transformers and CNNs, such as TransUnet [7], HiFormer [8] and LeVit-Unet [9]. These architectures combine the locality of CNNs with the long-range relevance of Transformers to encode both global and local features in medical image segmentation. However, although these methods demonstrate better performance in segmentation tasks, they still have limitations, such as insufficient attention to fine-grained features. The shortcoming hinders the model’s ability to accurately identify small targets and locate boundaries, thereby limiting its effectiveness in learning and generalizing in medical image segmentation. As illustrated in Figure 1, the PraNet model is used as an example to highlight the challenges in polyp segmentation.
Figure 1: Examples of segmentation of the PraNet model for different challenge cases.
Figure 2: The overall architecture of HiFSeg network. The entire model contains three components: (a) pyramid vision transformer (PVT) as encoder; (b) pyramid global-local interaction module(GLIM) for fusing multi-level features; (c) selective aggregation module(SAM) for integrating the high- and low-level features selectively for the final output.
Inspired by multiscale and multilevel feature modeling approaches [10, 11, 12, 13], we propose a high-frequency information enhanced polyp segmentation called HiFiSeg. The main components of HiFiSeg are pyramid vision transformer(PVT), global-local interaction module(GLIM), and selective aggregation module(SAM). PVT, a lightweight Transformer with a hierarchical structure, serves as the encoder to capture multiscale features. GLIM employs parallel convolutional kernels and pooling operations of various sizes to aggregate global and local information, enabling the extraction of fine-grained features, which is particularly beneficial for localizing small targets. To further reduce computational complexity, GLIM groups the channels and performs parallel convolutions using depthwise separable convolution. SAM enhances boundary features by using high-level semantic information to selectively guide low-level detail refinement. In summary, our contributions in this paper are as follows:
•We propose a novel framework for colon polyp segmentation, named HiFiSeg. The HiFiSeg adapts the pyramid vision transformer as an encoder to extract more robust features than CNN-based methods.
•We design global-local interaction module(GLIM) and selective aggregation module(SAM) to support our framework. GLIM enhances segmentation performance for small targets by extracting high-level local multi-scale information. SAM addresses boundary ambiguity by selectively fusing low-level boundary details with high-level semantic information.
•We conduct a set of experiments on five standard benchmark datasets for polyp segmentation(Kvasir [14], CVC-ClinicDB [15], CVC-300 [16], CVC-ColonDB [17], ETIS [18]) and compare the results with those from state-of-the-art methods.On the challenging CVC-ColonDB and ETIS datasets, the mDice scores achieved 0.826 and 0.822, respectively, surpassing existing state-of-the-art methods.
II Related Work
II-A Convolutional Neural Networks
CNNs are deep learning models specifically designed for processing image data, excelling in feature extraction capabilities, and are widely used in computer vision tasks. In recent years, CNN-based structures represented by the UNet architecture have made significant progress in medical image segmentation. UNet consists of a symmetric encoder and decoder, with skip connections that transfer features from the encoder to the decoder, combining low-level features and high-level semantic information to achieve high-precision segmentation. Many works have made improvements based on the UNet architecture, such as UNet++ [19], ResUNet++ [20] and DoubleUnet  [21].
Unlike the UNet-based methods, PolypNet [22] proposed a dual-tree wavelet pooling CNN with a local gradient-weighted embedding level set, significantly reducing the false positive rate, significantly reducing the false positive rate. Caranet [23] proposed a context axial reserve attention network to improve the segmentation performance on small objects. PraNet [24] generates a global map based on high-level features aggregated by the parallel partial decoder and employs the reverse attention module to mine boundary cues, effectively correcting any misaligned predictions, thereby improving segmentation accuracy.
Figure 3: Details of the introduced global-local interaction module(GLIM).It consists of three convolutional branches and a global average pooling branch.
II-B Vision Transformer
Transformer [5] is a deep learning model proposed by Vaswani et al. that uses a multi-head self-attention mechanism to capture long-range dependencies in input sequences. The Transformer model was initially applied to natural language processing such as machine translation and text generation. Due to its powerful parallel processing and ability to capture long-range dependencies, it is now widely used in image processing, speech recognition, and other fields. The Vision Transformer (ViT) [6], proposed by Dosovitskiy et al., is the first pure Transformer for image classification tasks. ViT divides an image into fixed-size patches, which are then projected and input into the Transformer model for processing and feature learning. Subsequently, Transformer-based pyramid structures applicable to downstream tasks have been proposed, such as Swin Transformer [25], PVT [10], and Segformer [26].
In medical image segmentation, Transfuse [27] combines Transformers and CNNs in a parallel fashion to efficiently capture global dependencies and low-level spatial details in a much shallower way. TransUNet [7] uses Transformer as encoder for medical image segmentation tasks, combined with U-Net to enhance local details by recovering local spatial information. Polyp-PVT [28] uses PVT as an encoder and designs a similarity aggregation module based on graph convolution. ColonFormer [29]models global semantic relations at multiple scales based on transformer and incorporates a new jump-joining technique for refining the boundaries of polyp objects in global maps. DuAT [30] proposes a dual-aggregate Transformer network in which the global-to-local spatial aggregation module is used to locate large and small targets, and the selective boundary aggregation module is used to preserve boundary details and locate recalibrated objects.
III Proposed Method
III-A Overview
As shown in Figure 2, our proposed network HiFiSeg consists of a pyramid vision transformer (PVT) encoder, global-local interaction module(GLIM), and selective aggregation module(SAM). The PVT encoder is employed to extract multi-scale hierarchical features from the input image $X\in\mathbb{R}^{H\times W\times 3}$, capturing both fine-grained local details and broad semantic information. Specifically, the PVT backbone yields four pyramid features $X_{i}\in\mathbb{R}^{\frac{H}{2^{i+1}}\times\frac{W}{2^{i+1}}\times C_{i}},\text{ where }i\in\{1,2,3,4\}\text{ and }C_{i}\in\{64,128,320,512\}$. The high-level features ${\{X_{i}|i\in\ (2,3,4)\}}$ are fed into the GLIM module to extract local multiscale features. The outputs of the GLIM module are then concatenated to produce the fused global-local multiscale feature $Output_{GLIM}$. The low-level feature $X_{1}$ is selectively aggregated with the high-level feature $X_{4}$ through the SAM module to obtain the enhanced edge feature $Output_{SAM}$. Finally, $Output_{GLIM}$ and $Output_{SAM}$ are fed into the segmentation heads to obtain the predicted results $P_{1}$ and $P_{2}$, respectively.
III-B Transformer Encoder
Some recent studies have demonstrated that pyramid structures, through the integration of multi-scale contextual information, can substantially improve the accuracy and efficiency of image segmentation. Our model uses the pyramid vision transformer (PVT) proposed in [46] as the encoder backbone to extract more robust features for polyp segmentation. PVT is the first pure Transformer backbone designed for various pixel-level dense prediction tasks. In polyp segmentation, PVT generates four multi-scale feature maps ${\{X_{i}|i\in\ (1,2,3,4)\}}$. Among these feature maps, $X_{1}$ gives detailed information about the polyps, while $X_{2},X_{3}$, and $X_{4}$ provide high-level features.
Figure 4: Qualitative results comparison of different models.
III-C Global-Local Interaction Module
In medical image segmentation, the context and background often occupy a much larger area than the segmentation target itself. Consequently, capturing information across different scales is essential for accurately segmenting small targets. Instead of presenting multi-scale features in a layer-by-layer fashion, GLIM aggregates global and local features at a specific high-level, achieving multi-scale feature representation at a granular level, reducing errors in high-level features.
The detailed architecture of our propoed GLIM is depicted in Figure 3, consisting of three convolution branches and one pooling branch. To balance accuracy and computational resources, we evenly divide the channels into four groups, applying depthwise separable convolution for each. After splitting the input features ${\{X_{i}\mid i\in\{2,3,4\}\}}$ into four components ${X_{i1},X_{i2},X_{i3},X_{i4}}$, they are fed into feature generation units at different scales. Given the input feature $X_{i}$, this procedure can be formulated as:
$X_{i1}$ $,X_{i2},X_{i3},X_{i4}=Split(X)$ (1)
$X^{{}^{\prime}}_{i1}$ $=Conv_{1\times 1}(X_{i1})$
$X^{{}^{\prime}}_{i2}$ $=DWConv_{3\times 3}(X_{i2})$
$X^{{}^{\prime}}_{i3}$ $=DWConv_{5\times 5}(X_{i3})$
$X^{{}^{\prime}}_{i4}$ $=\text{Sigmoid}(\text{GAP}(X_{i4}))\odot(X_{i4}),$
where $Split(\cdot)$ denotes the channel separation operation, $Conv_{1\times 1}(\cdot)$ represents 1×1 convolution, $DWConv_{3\times 3}(\cdot)$ refers to 3×3 depth-wise convolution, $DWConv_{5\times 5}(\cdot)$ refers to 5×5 depth-wise convolution, GAP stands for global average pooling, Sigmoid refers to the Sigmoid activation function, and $\odot$ is the element-wise product. The convolution branchs employ kernels of varying sizes to extract features at different scales of the image, while the pooling branch uses global average pooling to aggregate global information. These feature maps ${\{X^{{}^{\prime}}_{i1},X^{{}^{\prime}}_{i2},X^{{}^{\prime}}_{i3},X^{{}^{\prime}}_{i4}}\}$ are then concatenated along the channel dimension, and a 1×1 convolution is applied to aggregate both global and local information, resulting in a rich feature representation $X^{{}^{\prime}}_{i}$.This process can be expressed as:
$X^{{}^{\prime}}_{i}=Conv_{1\times 1}(Concat([X^{{}^{\prime}}_{i1},X^{{}^{\prime}}_{i2},X^{{}^{\prime}}_{i3},X^{{}^{\prime}}_{i4}])),$ (2)
where $Concat(\cdot)$ denotes a concatenation operation, while $Conv_{1\times 1}(\cdot)$ refers to a 1×1convolution. To enhance feature selection, we apply the GELU activation function to the feature $X^{{}^{\prime}}_{i}$ to generate the attention feature map, and then modulate the input feature $X$ through element-wise multiplication. It can be formulated as:
$X^{{}^{\prime\prime}}_{i}=GELU(X^{{}^{\prime}}_{i})\odot X_{i},$ (3)
where GELU refers to the GELU activation function, and $\odot$ is the element-wise product.
III-D Selective Aggregation Module
Shallow features contain rich spatial information, while deep features contain more semantic information. The effective combination of these two is crucial for improving the accuracy of the model. In order to enhance the guidance of shallow detail features by deep semantic features, we propose the selective aggregation module (SAM), as shown in Figure 1(c). Unlike previous fusion methods that directly add the provided feature maps, SAM selectively aggregates the features. First, the shallow feature $\hat{X}_{1}$ and deep feature $\hat{X}_{4}$ are individually processed through 1×1 convolutions followed by sigmoid activations to produce the attention weight $\sigma$. The output of the Sigmoid function could be represented as:
$\sigma=Sigmoid(C_{1\times 1}(\hat{X}_{1})\oplus C_{1\times 1}(\hat{X}_{4})).$ (4)
If $\sigma$ is high, the model assigns greater trust to the shallow feature, and vice versa.The output of the SAM can be written as:
$Output_{SAM}=\sigma\hat{X}_{1}\oplus(1-\sigma)\hat{X}_{4}.$ (5)
III-E Loss Function
We use weighted binary cross-entropy(BCE) loss and the weighted intersection over union(IoU) loss for supervision. Our loss function can be formulated as Eqn. 6:
$L_{total}$ $=L({P_{1},G})+L({P_{2},G})$ (6)
$L({P_{1},G})$ $=\lambda_{1}L^{w}_{BCE}(P_{1},G)+\lambda_{2}L^{w}_{IoU}(P_{1},G)$
$L({P_{2},G})$ $=\lambda_{1}L^{w}_{BCE}(P_{2},G)+\lambda_{2}L^{w}_{IoU}(P_{2},G),$
where $P_{1}$, $P_{2}$ are the outputs and G is the ground truth, $\lambda_{1}$ and $\lambda_{2}$ are the weighting coefficients, $L^{w}_{BCE}(\cdot)$ and $L^{w}_{IoU}(\cdot)$ are the weighted BCE and weighted IoU.
IV Experiment and Analysis
To validate the proposed HiFiSeg method’s superiority, it is compared with multiple state-of-the-art approaches on five popular datasets for polyp segmentation, namely, Kvasir [14], CVC-ClinicDB [15], CVC-300 [16], CVC-ColonDB [17], ETIS [18].
TABLE I: Quantitative Evaluation of the Learning Ability of Our Proposed HiFiSeg and Other Methods on the Kvasir and CVC-ClinicDB Datasets.
Methods Kvasir CVC-ClinicDB
mDice mIoU MAE mDice mIoU MAE
U-Net [2] 0.818 0.746 0.055 0.823 0.755 0.055
UNet++ [19] 0.821 0.743 0.048 0.794 0.729 0.022
Pranet [24] 0.898 0.840 0.030 0.899 0.849 0.009
SANet [31] 0.904 0.847 0.028 0.916 0.859 0.012
TransUnet [7] 0.913 0.857 0.028 0.935 0.887 0.008
SSFormer [32] 0.926 0.874 0.023 0.927 0.876 0.009
Polyp-PVT [28] 0.917 0.864 0.023 0.937 0.889 0.006
ColonFormer-S [29] 0.927 0.877 0.021 0.932 0.883 0.008
HiFiSeg(Ours) 0.933 0.886 0.018 0.942 0.897 0.006
TABLE II: Quantitative Evaluation of the Generalization Ability of Our Proposed HiFiSeg and Other Methods on the CVC-300, CVC-ColonDB and ETIS Datasets.
Methods CVC-300 CVC-ColonDB ETIS
mDice mIoU MAE mDice mIoU MAE mDice mIoU MAE
UNet [2] 0.710 0.627 0.022 0.512 0.444 0.061 0.398 0.335 0.036
UNet++ [19] 0.707 0.624 0.018 0.794 0.729 0.022 0.401 0.344 0.035
Pranet [24] 0.851 0.797 0.010 0.712 0.640 0.043 0.628 0.567 0.031
SANet [31] 0.888 0.815 0.008 0.753 0.670 0.043 0.750 0.654 0.015
TransUnet [7] 0.893 0.660 0.009 0.781 0.699 0.036 0.731 0.824 0.021
SSFormer [32] 0.887 0.821 0.007 0.772 0.697 0.036 0.767 0.698 0.016
Polyp-PVT [28] 0.900 0.833 0.007 0.808 0.727 0.031 0.787 0.706 0.013
ColonFormer-S [29] 0.894 0.826 0.008 0.811 0.730 0.027 0.789 0.711 0.013
HiFiSeg(Ours) 0.905 0.839 0.006 0.826 0.749 0.028 0.822 0.743 0.012
IV-A Datasets
We used five challenging public datasets for the polyp segmentation task, including Kvasir [14], CVC-ClinicDB [15], CVC-300 [16], CVC-ColonDB [17], and ETIS [18], to validate the learning and generalization capabilities of our model. Details for each dataset are as follows:
Kvasir dataset: The dataset consists of 1000 images with different resolutions from 720 × 576 to 1920 × 1072 pixels.
CVC-ClinicDB dataset:The dataset contains 612 polyp images which are extracted from 29 different endoscopic video clips.The resolution of images is 384 x 288.
CVC-300 dataset:The dataset consists of 60 polyp images and the resolution of the images is 574 x 500.
CVC-ColonDB dataset:The dataset consists of 380 polyp images and the resolution of the images is 570 x 500.
ETIS dataset:The dataset consists of 196 polyp images and the resolution of the images is 1225 x 966.
IV-B Evaluation Metrics
We employ three widely-used met- rics in the field of medical image segmentation,i.e., mean Dice (mDice), mean IoU (mIoU) and mean absolute error (MAE) to evaluate the model performances. Mean Dice and IoU are widely utilized metrics that primarily focus on assessing the internal consistency of segmentation results. MAE, on the other hand, measures the pixel-level accuracy by calculating the average absolute error between the predicted and actual values.
IV-C Implementation Details
[…]
We implement the HiFiSeg using the PyTorch framework, utilizing an NVIDIA RTX 3090 GPU. To enhance the model’s robustness concerning varying image sizes, the training images are scaled by factors of 0.75, 1, and 1.25 [29], respectively, before being fed into the model for learning. PVT encoder uses the same parameters as pvt_v2_b2 [33]. The model is trained end-to-end using the AdamW [34] optimizer, with the learning rate and weight decay set to 1e-4. The batch size is configured to 16.
TABLE III: QUANTITATIVE RESULTS FOR ABLATION STUDIES.
Methods Kvasir CVC-ClinicDB CVC-300 CVC-ColonDB ETIS
mDice mIoU mDice mIoU mDice mIoU mDice mIoU mDice mIoU
Baseline 0.910 0.856 0.903 0.847 0.869 0.792 0.796 0.707 0.759 0.668
w/o GLIM 0.924 0.878 0.937 0.892 0.892 0.828 0.813 0.734 0.798 0.725
w/o SAM 0.918 0.871 0.924 0.879 0.896 0.833 0.798 0.721 0.801 0.729
w/o Conv 0.923 0.877 0.931 0.883 0.886 0.816 0.808 0.732 0.788 0.710
w/o GAP 0.928 0.878 0.934 0.884 0.899 0.833 0.811 0.726 0.781 0.704
HiFiSeg(GLIM+SAM) 0.933 0.886 0.942 0.897 0.905 0.839 0.826 0.749 0.822 0.743
Figure 5: Visualization of the ablation study results.
IV-D Comparison with State-of-the-art Methods
IV-D1 Learning Ability
We first evaluate the learning ability of the proposed model HiFiSeg on the training datasets Kvasir and ClinicDB. As shown in Table I, we compare our proposed HiFiSeg with recently published and classical models for polyp segmentation, including CNN-based models such as UNet [2], UNet++ [19], PraNet [24], and SANet [31], as well as Transformer-based models like TransUnet [7], SSFormer [32], Polyp-PVT [28], and ColonFormer [29]. These results demonstrate the effectiveness of our model in accurately segmenting polyps. Specifically, the HiFiSeg model has a mDice value of 0.933 and a mIoU value of 0.876 on Kvasir dataset, which are 0.6% and 0.9% higher than the best performing model, ColonFormer, respectively. For CVC-ClinicDB dataset, the HiFiSeg model has a mDice value of 0.942 and a mIoU value of 0.897, which are 0.6% and 0.8% higher than the best performing model, Polyp-PVT, respectively.
IV-D2 Generalization Capabilities
To further evaluate our model’s generalization performance, we test HiFiSeg on three unseen datasets: CVC-300, CVC-ColonDB, and ETIS. These datasets originate from different medical centers, each presenting unique challenges and characteristics. As seen in Table II, on three unseen datasets, our model outperforms peer models across all metrics, demonstrating strong generalization performance. On CVC-300 dataset, HiFiSeg achieves mDice of 0.905 and mIoU of 0.839, outperforming the second-best model, Polyp-PVT, by 0.5% and 0.6%, respectively. On CVC-ColonDB dataset, our model’s mDice and mIoU scores are 1.5% and 1.9% higher than those of ColonFormer, respectively. Moreover, HiFiSeg achieves mDice of 0.822 and mIoU of 0.743 on ETIS dataset, which are 3.3% and 3.2% higher than the second-best model ColonFormer.
IV-D3 Visual Results
Figure 4 presents the visualization results of our model alongside the comparison models, providing a qualitative assessment of their performance. As shown in Figure 4, our model produces significantly fewer incorrectly predicted pixels in the segmentation results compared to other models. It accurately identifies colonic tissues and polyps, efficiently captures the boundaries of tiny polyps and target objects, and maintains stable recognition and segmentation capabilities across various imaging conditions. As seen in the first three rows of Figure 4, HiFiSeg accurately captures the boundaries and fine details of the target object, whereas the other methods fail to clearly detect the boundaries. In rows 4 and 5, our method demonstrates superior ability in identifying small targets and produces more accurate segmentation predictions.
IV-E Ablation Studies and Analysis
We use PVTv2 as our baseline (Bas.) and evaluate module effectiveness by removing components from the complete GLIM. The training, testing, and hyperparameter settings are the same as mentioned in Sec. III-C. The results are shown in Table III.
Effectiveness of GLIM. To evaluate the effectiveness of GLIM, we trained a version of the model: ”HiFiSeg (w/o GLIM).”As shown in Table III, compared to the standard HiFiSeg network, the performance of HiFiSeg (w/o GLIM) is reduced across all five datasets. This is particularly noticeable on ETIS dataset, where the mDice drops from 0.822 to 0.798 and the mIoU decreases from 0.743 to 0.725. As shown in the visualization results in Figure 5, the HiFiSeg (w/o GLIM) model struggles to effectively distinguish between polyps and colon tissues and has difficulty accurately localizing targets, particularly small ones. In contrast, the HiFiSeg model, with the inclusion of the GLIM module, significantly improves the accuracy of target localization and small target detection due to the aggregation of local and global features.
Effectiveness of SAM. To evaluate the effectiveness of SAM, we trained a version of the model: ”HiFiSeg (w/o SAM).” As shown in Table III, compared to HiFiSeg(w/o SAM), HiFiSeg shows a substantial improvement in performance on all five datasets. Specifically, the mdice on CVC-ColonDB dataset is improved from 0.798 to 0.826 and the mIoU is mentioned from 0.721 to 0.749, both of which are improved by 2.8%. The visualization results in Figure 5 show that SAM enables more accurate boundary extraction by effectively combining local pixel information with global semantic cues.
Effectiveness of GLIM module components. The GLIM consists of three convolution branches and one global average pooling(GAP) branch. The convolutional branches extract local features at multiple scales using convolutional kernels of varying sizes, while the global average pooling branch captures global information by spatially averaging the entire feature map, thereby better capturing the overall semantic context. To verify the effectiveness of the convolutional branches, we removed it from GLIM, resulting in HiFiSeg (w/o Conv). As shown in Table III, compared to the original HiFiSeg, the performance of the modified model drops significantly due to the lack of rich local representations, particularly on the ETIS dataset, where mDice and mIoU decrease by 3.4% and 3.3%, respectively. To verify the effectiveness of GAP, it is replaced with a 7×7 convolution, resulting in the model HiFiSeg (w/o GAP). As shown in Table III, the lack of global semantic information leads to a performance decline across all datasets, particularly on the ETIS dataset, where mDice and mIoU drop by 4.1% and 3.9%, respectively.
V Conclusion
In this paper, we proposed HiFiSeg network to address the challenges in colon polyp image segmentation, such as fine-grained target localization and boundary feature enhancement. Specifically, the GLIM fused global and local features by extracting multi-scale features in parallel, which facilitated the localization of targets of varying sizes. The SAM selectively combined semantic features with detailed features to alleviate the issue of unclear boundaries, further enhancing performance. Experimental results on five representative colon polyp datasets demonstrated that the HiFiSeg algorithm possessed strong learning and generalization capabilities, outperforming other competing methods. In future work, we plan to explore lightweight architectures to reduce model complexity, thereby extending its applicability to a wider range of medical image segmentation tasks.